how do you do a problem like:
3/4z=-5 1/2

Answer by AnlytcPhil(1807)

3/4z = -5 1/2

First change -5 1/2 to -11/2

3/4z = -11/2

Clear of fractions by multiplying both sides by LCD = 4

4(3/4z) = 4(-11/2)

     3x = -22

Divide both sides by 3

   3x/3 = -22/3

      x = -22/3

Or, if you prefer a mixed fraction:

      x = -7 1/3

Edwin
AnlyctPhil@aol.com